Consider the following hypothesis test: H0: μ = 18 Ha: μ ≠ 18 A sample of 48 provided a sample mean x = 17 and a sample standard deviation s = 4.9. a. Compute the value of the test statistic (to three decimal places.) b. Use the t distribution table (Table 2 in Appendix B) to compute a range for the p-value. (to two decimal places) p-value is between is c. At α = .05, what is your conclusion? p-value is Selectgreater than or equal to 0.05, rejectgreater than 0.05, do not rejectless than or equal to 0.05, do not rejectless than 0.05, rejectequal to 0.05, do not rejectnot equal to 0.05, do not rejectItem 4 H0 d. What is the rejection rule using the critical value? Reject H0 if t is Selectgreater than or equal to -2.012greater than 2.012less than or equal to -2.012less than -2.012equal to 2.012not equal to -2.012Item 5 or t is Selectgreater than or equal to 2.012greater than -2.012less than or equal to 2.012less than -2.012equal to 2.012not equal to -2.012Item 6 What is your conclusion? t = ; Selectdo not rejectrejectItem 8 H0
Answer #1

Here

sample mean \bar{x} = 17

sample standard deviation s = 4.9

and sample size n = 48

The test statistic can be written as

t = \frac{\sqrt{n}(\bar{x} - 18)}{s} which under H0 follows a t distribution with n-1 df.

We reject H0 at 5% level of significance if |t | > t_{0.025,n-1} = t_{0.025,47}=2.012 or P-value < 0.05

Now,

The value of the test statistic t_{obs} = -1.41392

and critical value t_{critical} = \pm t_{0.025,n-1} = \pm t_{0.025,47} = \pm 2.012

and P-value

=P(|t_{n-1}|>t_{obs}) = P(|t_{47}|>-1.41392) = 2*P(t_{47}<-1.41392) = 2*0.0819878 = 0.163976

Since P-value > 0.05 and |t_{obs}| = 1.41392 \ngtr 2.012 , so we fail to reject H0 at 5% level of significance and we can conclude that the population mean is not significantly different from 18.